OpenStudy (anonymous):

Quick Question! The product of (a - b)(a - b) is a perfect square trinomial, correct?

OpenStudy (anonymous):

Sometimes, always, or never?

jimthompson5910 (jim_thompson5910):

hint: x*x = x^2 by definition

OpenStudy (anonymous):

So, I'm taking that as an always :P

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

replace every x with (a-b) to get (a-b)(a-b) = (a-b)^2

OpenStudy (anonymous):

haha thanks. I was overthinking that a bit.

jimthompson5910 (jim_thompson5910):

you're welcome